# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v1

@@ 5.2 Multivariate Gaussian Target Distribution
Figure 13: ESJD vs swap acceptance rate for parallel tempering multivariate Gaussian target in dimension 20.
We first give an example of the standard multivariate Gaussian distribution to verify the correctness of our implementation. Note that the standard multivariate Gaussian as a target distribution satisfies the necessary assumptions of Theorem 1 in [1]: our target is a product of individual component densities, and we temper the distributions by raising them to the power of the original density, and $\epsilon\searrow 0$ such that $\epsilon=d^{-1/2}\ell$ for some positive constant $\ell$. We find that the results are consistent with the theoretical figure, and that the optimal swap acceptance rate is around 0.234. Please refer to Figure 13 for an illustration. Although the iterative temperature construction algorithm described in Section 5.1.1 may be costly in general, this example illustrates that if the practitioner is willing and able to construct the inverse temperatures with a spacing given by a swap acceptance probability, constructing a ladder with a swap acceptance probability of about 0.234 may be a viable and optimal strategy in the presence of certain restrictions on the target density.
5.3 Multimodal Distributions
5.3.1 Rough Carpet Distribution
Figure 14: Parallel tempering histogram and traceplot of the cold chain first component in the carpet distribution after 100,000 iterations.
In this section, we use the same target distribution as described in Section 3.4.1. Here, we examine the simpler case with homogeneous scaling factors $f_{d}(\mathbf{x})=\prod_{i=1}^{n}f(x_{i})$. Our individual component density $f(x)$ is the same as Equation (1), and in the $f(x)$ definition we set $\mu_{1}=-10$, $\mu_{2}=0$, $\mu_{3}=10$. We provide an example histogram and traceplot after 100,000 iterations with a constructed ladder to demonstrate what mixing looks like. Unlike the standard random-walk Metropolis which cannot escape from the central mode in this setting of the target distribution, the values in the cold chain swap between modes considerably often due to the swapping of values between chains. Refer to Figure 14 to see the cold chain’s mixing in the first component (all components are i.i.d.).
Figure 15: ESJD vs swap acceptance rate for the parallel tempering rough carpet target density in $d=20,30$.
Table 8: Constructed swap acceptance rates that maximise ESJD for the rough carpet product target density.
Dimension $d$ 20 30
Optimal Acceptance Rate 0.07051 0.04025
[…]
We conduct this experiment in $d=20,30$ and report the results in Figure 15 and Table 8. In both experiments, the optimal swap acceptance rate is less than 0.1, which is a substantial discrepancy from the theoretical figure of 0.234. In fact, the observed optimal swap acceptance rate decreases with increasing dimensionality. This might be due to the awkward shape of the distribution. Since it has $3^{d}$ modes, the density at each mode cannot be that high, which might mean the target distribution of a hot chain would look almost uniform and thus hot chains might not be sampling the right values that need to be swapped to colder chains.
5.3.2 Three Mixture Distribution
In this section, we use the same target distribution as described in Section 3.4.3 Equation 2. We set weights $w_{1}=w_{2}=w_{3}=\frac{1}{3}$, means $\mu_{1}=(-15,0,0,\dots,0)$, $\mu_{2}=(0,0,\dots,0)$, $\mu_{3}=(15,0,0,\dots,0)$ and covariances $\Sigma_{1}=\Sigma_{2}=\Sigma_{3}=I_{d}$. We provide an example histogram and traceplot after 100,000 iterations with a constructed ladder to demonstrate what the cold chain mixing looks like in the first component in Figure 16.
Figure 16: Parallel tempering histogram and traceplot of the cold chain first component in the Three-Mixture distribution after 100,000 iterations. Note that the rightmost mode is slightly undersampled.
Figure 17: ESJD vs swap acceptance rate for the parallel tempering three-mixture target density in $d=20,30$.
Table 9: Constructed swap acceptance rates that maximise ESJD for the three-mixture target density.
Dimension $d$ 20 30
Optimal Acceptance Rate 0.2520 0.2369
[…]
We refer the reader to Figure 17 and Table 9 for the results. Unlike the previous rough carpet density, we find that for the three-mixture density, constructing a ladder with a spacing of 0.234 seems approximately optimal even in low dimensions. However, the target density here is not in the i.i.d. product form. As noted in the Section 3.4.3 experiments, taking the sum of three Gaussians as a target density does seem to allow the RWM optimal scaling results to hold despite being outside of the theoretical assumptions, and this may extend to the parallel tempering optimal scaling framework too.
@@ 6 Discussion
We have provided an extensive list of experiments examining optimal scaling and acceptance rates for the Random Walk Metropolis (RWM) algorithm with a variety of target and proposal distributions. We have also examined the optimal scaling and acceptance rates for parallel tempering. Overall, the 0.234 acceptance rate for RWM is surprisingly robust and generalizable. Even in lower dimensions, it appears roughly optimal for non-Gaussian proposals and various multimodal target distributions where the target density is a product of i.i.d. single-dimensional component densities such as the “rough carpet”, or even a weighted sum of Gaussian densities (a mixture of three Gaussians) which is not included in the theorem assumptions. This result for the Gaussian mixture distribution even appears to extend to a inhomogeneous scaling of each component in each Gaussian, highlighting the robustness of the 0.234 acceptance rate figure and offering a suggestion for further relaxation of the necessary assumptions.
That being said, there are cases where 0.234 does not seem as relevant. The main case is in very low dimensions; in all our experiments with varying target densities, the optimal acceptance rate strays from 0.234 below a certain threshold which varies for each target density. Furthermore, although it is known that the 0.234 acceptance rate is not optimal for the continuous hypercube, this is somewhat surprising at first glance since this target density is continuously differentiable everywhere except on its boundary, which has zero measure. Lastly, the inhomogeneous scaling of each component in multimodal distributions is a crucial consideration affecting this optimal acceptance rate in practice; although in the infinite-dimensional limit and for large enough dimensions at about $d=50$, the heterogeneous scaling factors should not and do not appear to affect the optimal acceptance rate, we find that in smaller dimensions, including inhomogeneous scaling causes the observed optimal acceptance rate to be strictly less than 0.234. Open questions remain about to what extent or why the inhomogeneous scaling factors may reduce the optimal acceptance rate in smaller dimensions. A possible argument to answer the latter question is that the law of large numbers for the expected value of the sampled scaling factors may not be applicable in smaller dimensions. To avoid this issue altogether, we suggest to practitioners that variables should have transformations applied to ensure they have similar scaling, i.e. we emphasize normalizing variables to have the 0.234 acceptance rate be approximately optimal.
@@
Next, we discuss the optimal swap acceptance rate figure for parallel tempering and compare it with the optimal acceptance rate for random-walk Metropolis. On one hand, the results for the multivariate Gaussian and three-mixture density examples show that the 0.234 swap acceptance rate figure may be optimal even in lower dimensions. Although costly, if the practitioner is willing and able to construct the inverse temperature ladder iteratively with a spacing given by a swap acceptance probability, the spacing dictated by a 0.234 swap acceptance probability may be optimal for the practitioner. However, the theoretical 0.234 figure may not always be applicable: the “rough carpet” product density which satisfies the restrictive i.i.d. product form for the target density shows that the practical optimal swap acceptance rate might be on a case-by-case basis. It is difficult to be sure why this is the case: we hypothesize that perhaps the reason why a swap acceptance rate of 0.234 is more optimal in this case is due to the greater concentration of density in the state space for the multivariate Normal and the three-mixture example compared to the rough carpet density, where density is spread out in $3^{d}$ modes.
@@
Not only that, constructing a temperature ladder iteratively with the “optimal spacing” is challenging and time-consuming. It requires the ability to draw samples from the target distribution, which is itself the main goal of the parallel tempering algorithm. Without a direct way to sample from the target, a practitioner could run an MCMC algorithm with the $\beta$ value being proposed for the new addition to the ladder, but this may be time-consuming with many attempts required until a suitable $\beta$ is found for each step of the ladder. Since this optimal spacing may not yield the optimal ESJD in practice for finite, lower-dimensional targets and with a finite number of runs, an alternative approach is to adaptively adjust the spacings of the temperature ladder during the algorithm’s runs based on its recent ESJD performance over a recent window as guidance to practitioners. This is known as adaptive parallel tempering [13]. Previous literature shows that adaptive parallel tempering algorithms can achieve substantial efficiency gains over a standard geometric temperature spacing [29] while still converging to the target distribution [13] under the assumptions that the target distribution is sufficiently regular and has tails decaying faster than exponentially. Adaptive parallel tempering may suffice to be “good enough” for most practitioners whilst avoiding the costly temperature construction algorithm.
@@
Lastly, while we have established that the 0.234 acceptance rate for RWM is still fairly good even in lower dimensions and on some target distributions that may not have the i.i.d. product form, we provide some precautions to the practitioner. The 0.234 acceptance rate is a good heuristic to tune the algorithm to, yet there are still many cases where it is not optimal in terms of efficiency, such as the continuous hypercube example in Section 3.3. Even if the 0.234 acceptance rate is in fact optimal, the acceptance rate alone does not necessarily guarantee good samples, as illustrated by Figure 6. What literature [23, 27] indicates is that another very important thing in practice beyond just the acceptance rate is defining a good proposal covariance matrix $\Sigma_{n}$ that is approximately proportional to the target covariance matrix $\Sigma_{\pi}$; doing so can significantly increase the asymptotic relative efficiency. Realistically, it may be impossible to know $\Sigma_{\pi}$ in advance or even provide a good estimate for it. In fact, understanding how we can extend this result to correlated targets is an ongoing challenge [30] and immediately relevant to the practitioner since real-world data tends to have correlations between variables. Practitioners may consider using the adaptive Metropolis algorithm [11] to dynamically update the algorithm’s proposal covariance matrix. This adaptive algorithm is very useful and yields significant speed benefits in low and high dimensions [7, 25], but an important potential pitfall to note is that adaptive methods are not guaranteed to converge to the correct target density; in many cases, they may fail to converge or converge to something completely different [24, 25].
@@
The authors did not receive any funding for this submitted work and have no relevant financial or non-financial competing interests. The code implementing the experiments and generating the figures used in this work, and the experiment results stored in JSON format are publicly available on GitHub at https://github.com/aidanmrli/montecarlo.

### v3

@@ 5.3 Results
Figure 8: ESJD vs. swap acceptance rate for the parallel tempering algorithm with the $\pi_{2},\pi_{4},\pi_{6}$ target distributions in dimension $d=30$. Red dotted line indicates a swap acceptance rate of 0.234.
Notably, the Rough Carpet target ($\pi_{4}$) shows a significant deviation. The optimal swap acceptance rate is consistently higher than 0.234 and increases away from the theoretical value with increasing dimension. This is surprising because the Rough Carpet target does satisfy the i.i.d. product form and other conditions of Assumption 4.1 by Atchadé et al. [1]. The failure of the 0.234 rule here could be because the extreme multimodality of the target could reduce the practical utility of the ”hot” chains. The target possesses $3^{d}$ modes, and its probability mass is spread thinly across this vast landscape. For a hot chain where the inverse temperature $\beta$ is close to 0, the tempered density becomes extremely flat and loses the slight peak that a unimodal target would have. Consequently, it fails to efficiently sample from the specific modal regions that are of interest to the colder chains; the information flow from hot to cold chains is broken. To swap successfully more often, the temperature steps must be made smaller, which forces the swap acceptance rate to be higher.
@@ 6 Discussion
We have presented an extensive empirical investigation into the generalizability of the 0.234 acceptance rate heuristic for both Random-Walk Metropolis (RWM) and Parallel Tempering (PT) algorithms, using a variety of target and proposal distributions. Our results demonstrate that the 0.234 rule’s utility extends beyond the classical i.i.d. product setting of Roberts et al. [29]. The heuristic remains approximately optimal for: (i) non-Gaussian proposals that satisfy a spherical symmetry (shell) condition, such as the Laplace and Uniform distributions; (ii) multimodal targets such as the “rough carpet” i.i.d product with $3^{d}$ modes, and non-product targets constructed as a sum of isotropic components, such as our Three-Mixture model; and (iii) even for highly correlated, non-linear targets like the Full and Hybrid Rosenbrock in sufficiently high dimensions ($d\geq 20$). These findings provide empirical validation for the generalized theory of Sherlock [38], which predicts that as long as the target’s geometry is sufficiently regular and isotropic in the high-dimensional limit, the 0.234 rule should hold.
That being said, there are cases where 0.234 does not seem as relevant. The main case is in very low dimensions; in all our experiments with varying target densities, the optimal acceptance rate strays from 0.234 below a certain threshold which varies for each target density. In particular, targets composed of independent, low-dimensional anisotropic blocks (Even Rosenbrock) or those that are globally anisotropic in low dimensions (2D Rosenbrock), have an optimal acceptance rate significantly below 0.234. In these cases, we hypothesize the high-dimensional averaging effect that stabilizes the target’s geometry has not occurred, and the sampler’s performance is dictated by the challenging local geometry. As the dimension increases, the law of large numbers mitigates the effect of any single scaling factor or local correlation, so the target becomes effectively less eccentric, complying with the theory of Sherlock [38], and the optimal acceptance rate recovers towards 0.234. On the other hand, the continuous hypercube and Neal’s funnel completely fail the 0.234 heuristic. Although the hypercube is continuously differentiable everywhere except on its boundary, which has zero measure, an intuitive argument is that as $d$ increases, the probability that the proposal has all the values inside decreases. As for the funnel, no single proposal scaling can be globally efficient, which is just a limitation of the RWM algorithm.
@@
Next, we discuss the optimal swap acceptance rate figure for parallel tempering and compare it with the optimal acceptance rate for random-walk Metropolis. On one hand, the results for the multivariate Gaussian and three-mixture density examples show that the 0.234 swap acceptance rate figure may be optimal even in lower dimensions. Although costly, if the practitioner is willing and able to construct the inverse temperature ladder iteratively with a spacing given by a swap acceptance probability, the spacing dictated by a 0.234 swap acceptance probability may be optimal for the practitioner. However, the theoretical 0.234 figure may not always be applicable: despite satisfying the i.i.d. conditions of the PT optimal scaling theory [1], the extreme multimodality ($3^{d}$ modes) of the “rough carpet” product density appears to render the hot chains ineffective. We hypothesize that because the probability mass is spread so thinly, the hot chains fail to sample from meaningful modal regions, breaking the flow of information to the cold chain. This forces the algorithm to use smaller temperature steps (and thus higher swap acceptance rates) to achieve mixing, representing a practical failure of the 0.234 PT rule.
@@
Not only that, constructing a temperature ladder iteratively with the “optimal spacing” is challenging and time-consuming. It requires the ability to draw samples from the target distribution, which is itself the main goal of the parallel tempering algorithm. Without a direct way to sample from the target, a practitioner could run an MCMC algorithm with the $\beta$ value being proposed for the new addition to the ladder, but this may be time-consuming with many attempts required until a suitable $\beta$ is found for each step of the ladder.
Since this optimal spacing may not yield the optimal ESJD in practice for finite, lower-dimensional targets and with a finite number of runs, an alternative approach is to adaptively adjust the spacings of the temperature ladder during the algorithm’s runs based on its recent ESJD performance over a recent window as guidance to practitioners. This is known as adaptive parallel tempering [19]. Previous literature shows that adaptive parallel tempering algorithms can achieve substantial efficiency gains over a standard geometric temperature spacing [41] while still converging to the target distribution [19] under the assumptions that the target distribution is sufficiently regular and has tails decaying faster than exponentially. Adaptive parallel tempering may suffice to be “good enough” for most practitioners whilst avoiding the costly temperature construction algorithm.
@@
Lastly, while we have established that the 0.234 acceptance rate for RWM is still fairly good even in lower dimensions and on some target distributions that may not have the i.i.d. product form, we provide some precautions to the practitioner. The 0.234 acceptance rate is a good heuristic to tune the algorithm to, yet there are still many cases where it is not optimal in terms of efficiency, such as the continuous hypercube example in Section 3.5. Even if the 0.234 acceptance rate is in fact optimal, the acceptance rate alone does not necessarily guarantee good samples, as illustrated by Figure 2. What literature [33, 37] indicates is that another very important thing in practice beyond just the acceptance rate is defining a good proposal covariance matrix $\Sigma_{d}$ that is approximately proportional to the target covariance matrix $\Sigma_{\pi}$; doing so can significantly increase the asymptotic relative efficiency. Realistically, it may be impossible to know $\Sigma_{\pi}$ in advance or even provide a good estimate for it. In fact, understanding how we can extend this result to correlated targets is an ongoing challenge [42] and immediately relevant to the practitioner since real-world data tends to have correlations between variables. Practitioners may consider using the adaptive Metropolis algorithm [15] to dynamically update the algorithm’s proposal covariance matrix. This adaptive algorithm is very useful and yields significant speed benefits in low and high dimensions [8, 35], but an important potential pitfall to note is that adaptive methods are not guaranteed to converge to the correct target density; in many cases, they may fail to converge or converge to something completely different [34, 35].
@@
The authors did not receive any funding for this submitted work and have no relevant financial or non-financial competing interests. The code implementing the experiments and generating the figures used in this work are publicly available on GitHub at https://github.com/aidanmrli/rwm-pt-pytorch.
During the preparation of this revision, the authors used generative AI models (Gemini 2.5 Pro Preview 05-06, Claude 4.0 Sonnet) to assist with coding and manuscript revision. Generative AI was utilized to help generate, debug, and optimize code for the simulations. The new generated code based on PyTorch allows for the use of Graphics Processing Units (GPUs) when running RWM and PT algorithms to significantly speed up simulations and run more simulations at once. For manuscript revision, Generative AI was used to improve phrasing and clarity on text written by the authors. The authors thoroughly reviewed and edited all AI-generated output and take full responsibility for the scientific integrity and final content of this publication.
Appendix A Plots for RWM simulations
Figure 9: ESJD vs. acceptance rate for the i.i.d. Gamma(3, 2) target distribution $\pi_{2}$ (Eq 10) under RWM with a Gaussian proposal in dimensions $d\in\{2,5,10,30,50,100\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 10: ESJD vs. acceptance rate for the standard multivariate Gaussian target distribution $\pi_{3}$ under RWM with a multivariate Laplace proposal $Q_{L}$ (Eq 11) in dimensions $d\in\{2,5,10,30,50,100\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 11: ESJD vs. acceptance rate for the standard multivariate Gaussian target distribution $\pi_{3}$ under RWM with a multivariate Uniform proposal $Q_{U}$ (Eq 11) in dimensions $d\in\{2,5,10,30,50,100\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 12: ESJD vs. acceptance rate for the i.i.d. rough carpet target distribution $\pi_{4}$ (Eq 12) under RWM with a Gaussian proposal in dimensions $d\in\{20,30,50\}$ from left to right. Red dotted line indicates an acceptance rate of 0.234.
Figure 13: ESJD vs. acceptance rate for the inhomogeneously scaled i.i.d. rough carpet target distribution $\pi_{5}$ (Eq 13) under RWM with a Gaussian proposal in dimensions $d\in\{20,30,50\}$ from left to right. Red dotted line indicates an acceptance rate of 0.234.
Figure 14: ESJD vs. acceptance rate for the three mixture target distribution $\pi_{6}$ (Eq 14) under RWM with a Gaussian proposal in dimensions $d\in\{20,30,50\}$ from left to right. Red dotted line indicates an acceptance rate of 0.234.
Figure 15: ESJD vs. acceptance rate for the inhomogeneously scaled three mixture target distribution $\pi_{7}$ (Eq 14) under RWM with a Gaussian proposal in dimensions $d\in\{20,30,50\}$ from left to right. Red dotted line indicates an acceptance rate of 0.234.
Figure 16: ESJD vs. acceptance rate for the Full Rosenbrock target $\pi_{8}$ (Eq 16) under RWM with a Gaussian proposal in dimensions $d\in\{5,10,20,30\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 17: ESJD vs. acceptance rate for the Even Rosenbrock target $\pi_{9}$ (Eq 17) under RWM with a Gaussian proposal in dimensions $d\in\{5,10,20,30\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 18: ESJD vs. acceptance rate for the Hybrid Rosenbrock target $\pi_{10}$ (Eq 18) under RWM with a Gaussian proposal in dimensions $d\in\{5,9,19,29\}$ from top-left to bottom-right. Red dotted line indicates an acceptance rate of 0.234.
Figure 19: Traceplot of the first three dimensions of the cold chain from one of the parallel tempering simulations on the rough carpet target distribution $\pi_{4}$ (Eq 12).
Figure 20: Traceplot of the first three dimensions of the cold chain from one of the parallel tempering simulations on the three-mixture target distribution $\pi_{6}$ (Eq 14).